OpenStudy (anonymous):

A certain bacteria will triple in 6 hours. If the final count is 8 times the original count, how much time has passed?

OpenStudy (anonymous):

I think some context would be helpful. What level math class is this for?

OpenStudy (anonymous):

The Math Book is Advanced Mathematical Concepts. The course is Analytic Trig & Geometry And the chapter for this question is called Exponential And Logarithmic Functions. I hope that helps. I have never used this site before.

OpenStudy (whpalmer4):

If we let \(P_0\) be the initial population, then\[P(t) = P_03^{(t/6)}\]\[P(0) = P_03^0 = P_0\]\[P(6) = P_0 3^{(6/6)} = P_0*3^1 = 3*P_0\] Now we need to solve \[P(t) = P_03^{(t/6)} = 8*P_0\] Divide through by \(P_0\) and take the log of both sides, then solve for \(t\). Remember that \(\log a^n = n \log a\).

OpenStudy (anonymous):

oh ok thank you

OpenStudy (whpalmer4):

Do you have an answer?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

What is it?